OpenStudy (chloelourdes):

traveling at full speed on a river, it takes 4 hours for a motorboat to travel 18 upstream and 21 miles back downstream. if the current's speed is 2 mph, what is the maximum speed of the boat in still water?

OpenStudy (phi):

where did you get this question ?

OpenStudy (sshayer):

assume it takes t hours to go upstream'. let speed of motorboat in still water=x speed of current =2 (x-2)t=18 \[t=\frac{ 18 }{ x-2 }\] ...(1) (x+2)(4-t)=21 \[(x+2)\left( 4-\frac{ 18 }{ x-2 } \right)=21\] \[(x+2)\left( \frac{ 4x-8-18 }{ x-2 } \right)=21\] \[(x+2)(4x-26)=21x-42\] \[4x^2-26x+8x-52-21x+42=0\] 4x^2-39x-10=0 solve for x